OpenStudy (anonymous):

2/5v+3/5=5/3v+4 please help!!

OpenStudy (anonymous):

solve for v

OpenStudy (anonymous):

First bring 3/5 to right side via subtraction. 2/5v=5/3v+(4-3/5) then multiply both sides by 5/2 to get rid of it on the left side. Then bring v terms together and get rid of the coefficient. You then get v=-19/51.

OpenStudy (anonymous):

ohhh ok. thank u!

OpenStudy (anonymous):

(2/5)v-(5/3)v = 4-(3/5) (6/15)v-(25/15)v = (20/5) - (3/5) (-19/15)v = (17/5) v = (17/5)*(-15/19) v = -51/19

OpenStudy (anonymous):

ugh! i put the wrong answer! dang it. thank u nkili

OpenStudy (anonymous):

What percent o 5.92 is 37? would that be 2.8%??